# Impact of 3D macro physics and nuclear physics on the $p$ nuclei in $\mathrm{O}$-$\mathrm{C}$ shell mergers

### v1

@@ Appendix C Results of varying the input nuclear reactions
${}^{74}\mathrm{Se}$ ${}^{75}\mathrm{Se}(\gamma,n)$ $-0.95$ $-0.10$ ${}^{130}\mathrm{Ba}$ ${}^{126}\mathrm{Ba}(\gamma,\alpha)$ $-0.19$ $-0.01$
${}^{78}\mathrm{Kr}$ ${}^{79}\mathrm{Kr}(\gamma,n)$ $-0.91$ $-0.28$ ${}^{132}\mathrm{Ce}(\gamma,n)$ $-0.23$ $-0.01$
${}^{84}\mathrm{Sr}$ ${}^{85}\mathrm{Sr}(\gamma,n)$ $-0.92$ $-0.26$ ${}^{132}\mathrm{Ce}(\gamma,p)$ $0.32$ $0.01$
${}^{92}\mathrm{Mo}$ ${}^{93}\mathrm{Mo}(\gamma,n)$ $-0.57$ $-0.03$ ${}^{132}\mathrm{Ce}(\gamma,\alpha)$ $-0.20$ $-0.01$
${}^{100}\mathrm{Pd}(\gamma,p)$ $0.19$ $0.01$ ${}^{134}\mathrm{Ce}(\gamma,n)$ $-0.20$ $-0.01$
@@
${}^{110}\mathrm{Sn}(\gamma,n)$ $0.16$ $0.01$ ${}^{132}\mathrm{Ba}$ ${}^{132}\mathrm{Ce}(\gamma,n)$ $-0.28$ $-0.04$
${}^{110}\mathrm{Sn}(\gamma,p)$ $0.23$ $0.01$ ${}^{132}\mathrm{Ce}(\gamma,p)$ $-0.25$ $-0.03$
${}^{110}\mathrm{Sn}(\gamma,\alpha)$ $0.37$ $0.02$ ${}^{132}\mathrm{Ce}(\gamma,\alpha)$ $-0.31$ $-0.04$
${}^{94}\mathrm{Mo}$ ${}^{93}\mathrm{Mo}(\gamma,n)$ $0.89$ $0.08$ ${}^{133}\mathrm{Ce}(\gamma,n)$ $-0.19$ $-0.03$
${}^{96}\mathrm{Ru}$ ${}^{97}\mathrm{Ru}(\gamma,n)$ $-0.62$ $-0.05$ ${}^{134}\mathrm{Ce}(\gamma,\alpha)$ $-0.18$ $-0.02$
${}^{97}\mathrm{Ru}(\gamma,\alpha)$ $-0.18$ $-0.01$ ${}^{138}\mathrm{La}$ ${}^{137}\mathrm{La}(\gamma,n)$ $-0.75$ $-0.45$
${}^{100}\mathrm{Pd}(\gamma,\alpha)$ $0.33$ $0.03$ ${}^{136}\mathrm{Ce}$ ${}^{138}\mathrm{Nd}(\gamma,n)$ $-0.42$ $-0.09$
@@
${}^{110}\mathrm{Sn}(\gamma,\alpha)$ $0.23$ $0.02$ ${}^{138}\mathrm{Ce}$ ${}^{138}\mathrm{Nd}(\gamma,n)$ $-0.33$ $-0.04$
${}^{98}\mathrm{Ru}$ ${}^{100}\mathrm{Pd}(\gamma,p)$ $0.70$ $0.13$ ${}^{138}\mathrm{Nd}(\gamma,p)$ $-0.31$ $-0.04$
${}^{100}\mathrm{Pd}(\gamma,\alpha)$ $-0.53$ $-0.10$ ${}^{138}\mathrm{Nd}(\gamma,\alpha)$ $-0.24$ $-0.02$
${}^{102}\mathrm{Pd}$ ${}^{100}\mathrm{Pd}(\gamma,p)$ $-0.39$ $-0.06$ ${}^{139}\mathrm{Nd}(\gamma,n)$ $-0.31$ $-0.04$
${}^{100}\mathrm{Pd}(\gamma,\alpha)$ $-0.38$ $-0.05$ ${}^{144}\mathrm{Sm}$ ${}^{142}\mathrm{Sm}(\gamma,n)$ $-0.36$ $-0.06$
${}^{103}\mathrm{Pd}(\gamma,n)$ $-0.23$ $-0.04$ ${}^{142}\mathrm{Sm}(\gamma,p)$ $-0.29$ $-0.04$
${}^{104}\mathrm{Cd}(\gamma,p)$ $0.17$ $0.03$ ${}^{143}\mathrm{Sm}(\gamma,n)$ $-0.50$ $-0.09$
${}^{104}\mathrm{Cd}(\gamma,\alpha)$ $-0.27$ $-0.04$ ${}^{196}\mathrm{Pb}(\gamma,n)$ $0.19$ $0.03$
${}^{106}\mathrm{Cd}$ ${}^{104}\mathrm{Cd}(\gamma,p)$ $-0.23$ $-0.03$ ${}^{152}\mathrm{Gd}$ ${}^{152}\mathrm{Dy}(\gamma,\alpha)$ $-0.50$ $-0.05$
${}^{107}\mathrm{Cd}(\gamma,n)$ $-0.62$ $-0.07$ ${}^{158}\mathrm{Er}(\gamma,n)$ $0.18$ $0.01$
${}^{110}\mathrm{Sn}(\gamma,p)$ $0.20$ $0.02$ ${}^{156}\mathrm{Dy}$ ${}^{156}\mathrm{Er}(\gamma,\alpha)$ $-0.44$ $-0.27$
@@
${}^{108}\mathrm{Cd}$ ${}^{107}\mathrm{Cd}(\gamma,n)$ $0.61$ $0.11$ ${}^{158}\mathrm{Er}(\gamma,\alpha)$ $-0.18$ $-0.07$
@@
${}^{110}\mathrm{Sn}(\gamma,\alpha)$ $-0.33$ $-0.06$ ${}^{158}\mathrm{Dy}$ ${}^{158}\mathrm{Er}(\gamma,n)$ $-0.28$ $-0.02$
${}^{113}\mathrm{In}$ ${}^{113}\mathrm{Sn}(\gamma,n)$ $0.94$ $0.44$ ${}^{158}\mathrm{Er}(\gamma,\alpha)$ $-0.53$ $-0.05$
${}^{112}\mathrm{Sn}$ ${}^{110}\mathrm{Sn}(\gamma,p)$ $-0.25$ $-0.03$ ${}^{162}\mathrm{Er}$ ${}^{162}\mathrm{Yb}(\gamma,\alpha)$ $-0.49$ $-0.30$
${}^{110}\mathrm{Sn}(\gamma,\alpha)$ $-0.34$ $-0.04$ ${}^{164}\mathrm{Yb}(\gamma,\alpha)$ $-0.16$ $-0.07$
${}^{113}\mathrm{Sn}(\gamma,n)$ $-0.65$ $-0.08$ ${}^{168}\mathrm{Hf}(\gamma,n)$ $0.17$ $0.07$
${}^{114}\mathrm{Sn}$ ${}^{110}\mathrm{Sn}(\gamma,\alpha)$ $-0.21$ $-0.02$ ${}^{164}\mathrm{Er}$ ${}^{164}\mathrm{Yb}(\gamma,n)$ $-0.31$ $-0.17$
${}^{113}\mathrm{Sn}(\gamma,n)$ $0.78$ $0.08$ ${}^{164}\mathrm{Yb}(\gamma,\alpha)$ $-0.56$ $-0.51$
${}^{122}\mathrm{Xe}(\gamma,n)$ $-0.28$ $-0.03$ ${}^{202}\mathrm{Pb}(\gamma,n)$ $0.16$ $0.09$
${}^{122}\mathrm{Xe}(\gamma,p)$ $0.16$ $0.02$ ${}^{168}\mathrm{Yb}$ ${}^{168}\mathrm{Hf}(\gamma,n)$ $-0.28$ $-0.12$
${}^{122}\mathrm{Xe}(\gamma,\alpha)$ $0.23$ $0.03$ ${}^{168}\mathrm{Hf}(\gamma,\alpha)$ $-0.55$ $-0.57$
${}^{169}\mathrm{Lu}(\gamma,n)$ $0.15$ $0.01$ ${}^{174}\mathrm{Hf}$ ${}^{174}\mathrm{W}(\gamma,n)$ $-0.30$ $-0.13$
${}^{115}\mathrm{Sn}$ ${}^{110}\mathrm{Sn}(\gamma,\alpha)$ $-0.21$ $-0.02$ ${}^{174}\mathrm{W}(\gamma,\alpha)$ $-0.53$ $-0.37$
${}^{113}\mathrm{Sn}(\gamma,n)$ $0.80$ $0.08$ ${}^{180}\mathrm{Ta}$ ${}^{179}\mathrm{Ta}(\gamma,n)$ $-0.87$ $-0.00$
${}^{122}\mathrm{Xe}(\gamma,n)$ $-0.27$ $-0.03$ ${}^{180}\mathrm{W}$ ${}^{180}\mathrm{Os}(\gamma,n)$ $-0.38$ $-0.21$
${}^{122}\mathrm{Xe}(\gamma,p)$ $0.15$ $0.02$ ${}^{180}\mathrm{Os}(\gamma,\alpha)$ $-0.52$ $-0.46$
@@
${}^{169}\mathrm{Lu}(\gamma,n)$ $0.15$ $0.01$ ${}^{184}\mathrm{Os}$ ${}^{184}\mathrm{Pt}(\gamma,n)$ $-0.16$ $-0.05$
${}^{120}\mathrm{Te}$ ${}^{120}\mathrm{Xe}(\gamma,\alpha)$ $-0.22$ $-0.05$ ${}^{184}\mathrm{Pt}(\gamma,\alpha)$ $-0.55$ $-0.29$
@@
${}^{122}\mathrm{Xe}(\gamma,\alpha)$ $-0.49$ $-0.10$ ${}^{190}\mathrm{Pt}$ ${}^{190}\mathrm{Hg}(\gamma,n)$ $-0.30$ $-0.13$
${}^{124}\mathrm{Xe}$ ${}^{122}\mathrm{Xe}(\gamma,n)$ $-0.28$ $-0.10$ ${}^{190}\mathrm{Hg}(\gamma,\alpha)$ $-0.49$ $-0.30$
${}^{122}\mathrm{Xe}(\gamma,p)$ $-0.31$ $-0.11$ ${}^{196}\mathrm{Pb}(\gamma,n)$ $0.20$ $0.08$
${}^{122}\mathrm{Xe}(\gamma,\alpha)$ $-0.43$ $-0.21$ ${}^{196}\mathrm{Hg}$ ${}^{196}\mathrm{Pb}(\gamma,n)$ $-0.76$ $-0.60$
@@
${}^{126}\mathrm{Xe}$ ${}^{126}\mathrm{Ba}(\gamma,p)$ $-0.35$ $-0.17$
${}^{126}\mathrm{Ba}(\gamma,\alpha)$ $-0.48$ $-0.32$
${}^{127}\mathrm{Ba}(\gamma,n)$ $-0.22$ $-0.12$

### v2

@@ Appendix C Results of varying the input nuclear reactions
$^{74}\mathrm{Se}$ $^{75}\mathrm{Se}(\gamma,n)$ 0.95-0.95 0.10-0.10 $^{130}\mathrm{Ba}$ $^{126}\mathrm{Ba}(\gamma,\alpha)$ 0.19-0.19 0.01-0.01
$^{78}\mathrm{Kr}$ $^{79}\mathrm{Kr}(\gamma,n)$ 0.91-0.91 0.28-0.28 $^{132}\mathrm{Ce}(\gamma,n)$ 0.23-0.23 0.01-0.01
$^{84}\mathrm{Sr}$ $^{85}\mathrm{Sr}(\gamma,n)$ 0.92-0.92 0.26-0.26 $^{132}\mathrm{Ce}(\gamma,p)$ $0.32$ $0.01$
$^{92}\mathrm{Mo}$ $^{93}\mathrm{Mo}(\gamma,n)$ 0.57-0.57 0.03-0.03 $^{132}\mathrm{Ce}(\gamma,\alpha)$ 0.20-0.20 0.01-0.01
$^{100}\mathrm{Pd}(\gamma,p)$ $0.19$ $0.01$ $^{134}\mathrm{Ce}(\gamma,n)$ 0.20-0.20 0.01-0.01
@@
$^{110}\mathrm{Sn}(\gamma,n)$ $0.16$ $0.01$ $^{132}\mathrm{Ba}$ $^{132}\mathrm{Ce}(\gamma,n)$ 0.28-0.28 0.04-0.04
$^{110}\mathrm{Sn}(\gamma,p)$ $0.23$ $0.01$ $^{132}\mathrm{Ce}(\gamma,p)$ 0.25-0.25 0.03-0.03
$^{110}\mathrm{Sn}(\gamma,\alpha)$ $0.37$ $0.02$ $^{132}\mathrm{Ce}(\gamma,\alpha)$ 0.31-0.31 0.04-0.04
$^{94}\mathrm{Mo}$ $^{93}\mathrm{Mo}(\gamma,n)$ $0.89$ $0.08$ $^{133}\mathrm{Ce}(\gamma,n)$ 0.19-0.19 0.03-0.03
$^{96}\mathrm{Ru}$ $^{97}\mathrm{Ru}(\gamma,n)$ 0.62-0.62 0.05-0.05 $^{134}\mathrm{Ce}(\gamma,\alpha)$ 0.18-0.18 0.02-0.02
$^{97}\mathrm{Ru}(\gamma,\alpha)$ 0.18-0.18 0.01-0.01 $^{138}\mathrm{La}$ $^{137}\mathrm{La}(\gamma,n)$ 0.75-0.75 0.45-0.45
$^{100}\mathrm{Pd}(\gamma,\alpha)$ $0.33$ $0.03$ $^{136}\mathrm{Ce}$ $^{138}\mathrm{Nd}(\gamma,n)$ 0.42-0.42 0.09-0.09
@@
$^{110}\mathrm{Sn}(\gamma,\alpha)$ $0.23$ $0.02$ $^{138}\mathrm{Ce}$ $^{138}\mathrm{Nd}(\gamma,n)$ 0.33-0.33 0.04-0.04
$^{98}\mathrm{Ru}$ $^{100}\mathrm{Pd}(\gamma,p)$ $0.70$ $0.13$ $^{138}\mathrm{Nd}(\gamma,p)$ 0.31-0.31 0.04-0.04
$^{100}\mathrm{Pd}(\gamma,\alpha)$ 0.53-0.53 0.10-0.10 $^{138}\mathrm{Nd}(\gamma,\alpha)$ 0.24-0.24 0.02-0.02
$^{102}\mathrm{Pd}$ $^{100}\mathrm{Pd}(\gamma,p)$ 0.39-0.39 0.06-0.06 $^{139}\mathrm{Nd}(\gamma,n)$ 0.31-0.31 0.04-0.04
$^{100}\mathrm{Pd}(\gamma,\alpha)$ 0.38-0.38 0.05-0.05 $^{144}\mathrm{Sm}$ $^{142}\mathrm{Sm}(\gamma,n)$ 0.36-0.36 0.06-0.06
$^{103}\mathrm{Pd}(\gamma,n)$ 0.23-0.23 0.04-0.04 $^{142}\mathrm{Sm}(\gamma,p)$ 0.29-0.29 0.04-0.04
$^{104}\mathrm{Cd}(\gamma,p)$ $0.17$ $0.03$ $^{143}\mathrm{Sm}(\gamma,n)$ 0.50-0.50 0.09-0.09
$^{104}\mathrm{Cd}(\gamma,\alpha)$ 0.27-0.27 0.04-0.04 $^{196}\mathrm{Pb}(\gamma,n)$ $0.19$ $0.03$
$^{106}\mathrm{Cd}$ $^{104}\mathrm{Cd}(\gamma,p)$ 0.23-0.23 0.03-0.03 $^{152}\mathrm{Gd}$ $^{152}\mathrm{Dy}(\gamma,\alpha)$ 0.50-0.50 0.05-0.05
$^{107}\mathrm{Cd}(\gamma,n)$ 0.62-0.62 0.07-0.07 $^{158}\mathrm{Er}(\gamma,n)$ $0.18$ $0.01$
$^{110}\mathrm{Sn}(\gamma,p)$ $0.20$ $0.02$ $^{156}\mathrm{Dy}$ $^{156}\mathrm{Er}(\gamma,\alpha)$ 0.44-0.44 0.27-0.27
@@
$^{108}\mathrm{Cd}$ $^{107}\mathrm{Cd}(\gamma,n)$ $0.61$ $0.11$ $^{158}\mathrm{Er}(\gamma,\alpha)$ 0.18-0.18 0.07-0.07
@@
$^{110}\mathrm{Sn}(\gamma,\alpha)$ 0.33-0.33 0.06-0.06 $^{158}\mathrm{Dy}$ $^{158}\mathrm{Er}(\gamma,n)$ 0.28-0.28 0.02-0.02
$^{113}\mathrm{In}$ $^{113}\mathrm{Sn}(\gamma,n)$ $0.94$ $0.44$ $^{158}\mathrm{Er}(\gamma,\alpha)$ 0.53-0.53 0.05-0.05
$^{112}\mathrm{Sn}$ $^{110}\mathrm{Sn}(\gamma,p)$ 0.25-0.25 0.03-0.03 $^{162}\mathrm{Er}$ $^{162}\mathrm{Yb}(\gamma,\alpha)$ 0.49-0.49 0.30-0.30
$^{110}\mathrm{Sn}(\gamma,\alpha)$ 0.34-0.34 0.04-0.04 $^{164}\mathrm{Yb}(\gamma,\alpha)$ 0.16-0.16 0.07-0.07
$^{113}\mathrm{Sn}(\gamma,n)$ 0.65-0.65 0.08-0.08 $^{168}\mathrm{Hf}(\gamma,n)$ $0.17$ $0.07$
$^{114}\mathrm{Sn}$ $^{110}\mathrm{Sn}(\gamma,\alpha)$ 0.21-0.21 0.02-0.02 $^{164}\mathrm{Er}$ $^{164}\mathrm{Yb}(\gamma,n)$ 0.31-0.31 0.17-0.17
$^{113}\mathrm{Sn}(\gamma,n)$ $0.78$ $0.08$ $^{164}\mathrm{Yb}(\gamma,\alpha)$ 0.56-0.56 0.51-0.51
$^{122}\mathrm{Xe}(\gamma,n)$ 0.28-0.28 0.03-0.03 $^{202}\mathrm{Pb}(\gamma,n)$ $0.16$ $0.09$
$^{122}\mathrm{Xe}(\gamma,p)$ $0.16$ $0.02$ $^{168}\mathrm{Yb}$ $^{168}\mathrm{Hf}(\gamma,n)$ 0.28-0.28 0.12-0.12
$^{122}\mathrm{Xe}(\gamma,\alpha)$ $0.23$ $0.03$ $^{168}\mathrm{Hf}(\gamma,\alpha)$ 0.55-0.55 0.57-0.57
$^{169}\mathrm{Lu}(\gamma,n)$ $0.15$ $0.01$ $^{174}\mathrm{Hf}$ $^{174}\mathrm{W}(\gamma,n)$ 0.30-0.30 0.13-0.13
$^{115}\mathrm{Sn}$ $^{110}\mathrm{Sn}(\gamma,\alpha)$ 0.21-0.21 0.02-0.02 $^{174}\mathrm{W}(\gamma,\alpha)$ 0.53-0.53 0.37-0.37
$^{113}\mathrm{Sn}(\gamma,n)$ $0.80$ $0.08$ $^{180}\mathrm{Ta}$ $^{179}\mathrm{Ta}(\gamma,n)$ 0.87-0.87 0.00-0.00
$^{122}\mathrm{Xe}(\gamma,n)$ 0.27-0.27 0.03-0.03 $^{180}\mathrm{W}$ $^{180}\mathrm{Os}(\gamma,n)$ 0.38-0.38 0.21-0.21
$^{122}\mathrm{Xe}(\gamma,p)$ $0.15$ $0.02$ $^{180}\mathrm{Os}(\gamma,\alpha)$ 0.52-0.52 0.46-0.46
@@
$^{169}\mathrm{Lu}(\gamma,n)$ $0.15$ $0.01$ $^{184}\mathrm{Os}$ $^{184}\mathrm{Pt}(\gamma,n)$ 0.16-0.16 0.05-0.05
$^{120}\mathrm{Te}$ $^{120}\mathrm{Xe}(\gamma,\alpha)$ 0.22-0.22 0.05-0.05 $^{184}\mathrm{Pt}(\gamma,\alpha)$ 0.55-0.55 0.29-0.29
@@
$^{122}\mathrm{Xe}(\gamma,\alpha)$ 0.49-0.49 0.10-0.10 $^{190}\mathrm{Pt}$ $^{190}\mathrm{Hg}(\gamma,n)$ 0.30-0.30 0.13-0.13
$^{124}\mathrm{Xe}$ $^{122}\mathrm{Xe}(\gamma,n)$ 0.28-0.28 0.10-0.10 $^{190}\mathrm{Hg}(\gamma,\alpha)$ 0.49-0.49 0.30-0.30
$^{122}\mathrm{Xe}(\gamma,p)$ 0.31-0.31 0.11-0.11 $^{196}\mathrm{Pb}(\gamma,n)$ $0.20$ $0.08$
$^{122}\mathrm{Xe}(\gamma,\alpha)$ 0.43-0.43 0.21-0.21 $^{196}\mathrm{Hg}$ $^{196}\mathrm{Pb}(\gamma,n)$ 0.76-0.76 0.60-0.60
@@
$^{126}\mathrm{Xe}$ $^{126}\mathrm{Ba}(\gamma,p)$ 0.35-0.35 0.17-0.17
$^{126}\mathrm{Ba}(\gamma,\alpha)$ 0.48-0.48 0.32-0.32
$^{127}\mathrm{Ba}(\gamma,n)$ 0.22-0.22 0.12-0.12